# Learning Stackable and Skippable LEGO Bricks for Efficient, Reconfigurable, and Variable-Resolution Diffusion Modeling

### v1

@@ 4.2 Main results
LEGO-XL-PR (ours, $512\times 512$) 9.01 4.18 6.39 4.47 168.92 227.06 0.76 0.85 0.63 0.50 512
Training in pixel space: We begin by comparing LEGO in pixel-space training on CelebA and ImageNet with two Transformer-based state-of-the-art diffusion models (Bao et al., 2022; Peebles & Xie, 2022), which are our most relevant baselines. All images are resized to a resolution of $64\times 64$. We conduct experiments with two spatial refinement methods using LEGO, denoted as -PG for the progressive growth variant and -PR for the progressive refinement variant, as explained in Section 3.3.2. In both Tables 2 and 2, when compared to DiT, LEGO slightly increases the number of parameters due to the expansion of the input dimension in a channel-wise manner. However, in terms of FLOPs, we observe that LEGO incurs a significantly lower computational cost compared to both U-ViT and DiT. Despite maintaining such a low computational cost, LEGO still outperforms the competition in terms of FID performance.
@@
Training in latent space: We conducted comparisons with state-of-the-art class-conditional generative models trained on ImageNet at resolutions of $256\times 256$ and $512\times 512$, as shown in Table 3. Our results were obtained after training with 512M images (calculated as the number of iterations $\times$ batch size). When considering LEGO without classifier-free guidance (CFG), we observe that it achieves a better FID than the baseline diffusion models, striking a good balance between precision and recall. One potential explanation for this improvement is that the LEGO bricks with small patch sizes enhance patch diversity, preserving both generation fidelity and diversity.
@@
With the inclusion of CFG, we notably achieved an FID of 2.05 and the best IS of 289.12 on ImageNet-$256\times 256$. In the case of using $512\times 512$ resolution data, LEGO also achieved the best IS and produced a competitive FID with fewer iterated images. Additionally, as depicted in Figure 1, LEGO demonstrated superior convergence speed and required significantly less training time. For qualitative validation, we provide visualizations of randomly generated samples in Figure 2.
4.3 Additional results and ablation studies
In Appendices B and C, we provide additional results and delve into ablation studies aimed at investigating the allocation of diffusion timesteps to LEGO bricks operating on different patch sizes and their corresponding impacts on performance. We illustrate in Appendix C.3 the versatility of LEGO by incorporating a pretrained diffusion model as the first LEGO brick of the progressive-refinement construction, which is either frozen or finetuned during training.
@@ Acknowledgements
H. Zheng, Z. Wang, and M. Zhou acknowledge the support of NSF-IIS 2212418, NIH-R37 CA271186, the Fall 2022 McCombs REG award, and the NSF AI Institute for Foundations of Machine Learning (IFML).
Reproducibility Statement
To ensure the reproducibility of our work, we include experimental details in Appendix B. PyTorch code for facilitating reproducible research will be made available on GitHub.
Appendix A Limitations and Future Work
@@
The current work exhibits several limitations. Firstly, we have not explored the application of LEGO bricks for text-guided image generation. Secondly, in generating class-conditional images at resolutions much higher than the training data, categories with low diversity may produce results that lack photorealism. Thirdly, we have primarily focused on stacking LEGO bricks in either a progressive-growth or progressive-refinement manner, leaving other stacking strategies unexplored. Fourthly, our selection of which LEGO bricks to skip at different time steps during generation relies on heuristic criteria, and there is room for developing more principled skipping strategies.
@@
Appendix B Main Experiments: Additional Details
Dataset and Training details
@@ Appendix B Main Experiments: Additional Details
[…]
@@
We provide an overview of the architecture configurations for the LEGO models with different capacities in Tables 4 and 5. Since Table 5 only differs from Table 4 in the brick sizes and local receptive field sizes, we will focus on explaining Table 4 in detail. Table 4 consists of three LEGO bricks, with two of them being patch-level bricks and one being an image-level brick. When using the PG-based spatial refinement approach, we stack the bricks from the bottom to the top following the order: $4\times 4\rightarrow 16\times 16\rightarrow 64\times 64$. In other words, the $4\times 4$ patch-level brick is placed at the bottom, followed by the $16\times 16$ patch-level brick, and then the $64\times 64$ image-level brick. Conversely, for the PR-based spatial refinement approach, we stack the bricks from the bottom to the top following the order: $64\times 64\rightarrow 16\times 16\rightarrow 4\times 4$. Adhering to conventional modeling techniques, we’ve outlined three distinct configurations — Small (LEGO-S), Large (LEGO-L), and XLarge (LEGO-XL) — used in our main experiments to cater to varying model capacities. Across bricks, the main differences are in token embedding and the number of attention heads, which revolves around the feature size of a local receptive field. Each brick specifies its patch size $r_{k}\times r_{k}$, local receptive field size $\ell_{k}\times\ell_{k}$, and embedding dimension (number of channels) $d_{k}$. The number of tokens (attention span) is determined as $(r_{k}/\ell_{k})^{2}$. In each brick, the configurations for LEGO-S, LEGO-L, and LEGO-XL also differ in terms of the number of attention heads and the number of DiT blocks used in the current brick to ensure scalability and adaptability across the architecture. To maintain generation coherence, we set the local-receptive-field size of the image-level brick as $2\times 2$, which effectively turns it into a DiT with a long attention span, albeit with significantly fewer layers compared to standard DiT models. Moreover, the image-level brick can be safely skipped at appropriate time steps during generation. We find the proposed configuration strikes a favorable balance between computational efficiency and performance. Please refer to our ablation studies on the effects of LEGO bricks in Appendix C.2. In training, LEGO patch-brick can operate on sampled patches to further save memory and computation. In our experiments on CelebA and ImageNet datasets, 50% and 75% of all non-overlapping patches are used in the training of patch-bricks, respectively. As noted in Section 3.2, there are instances where $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ contains missing values as some patches are not sampled in the preceding brick. In such cases, we replace these missing values with the corresponding pixels from ${\mathbf{x}}_{0}$ to ensure that the required $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ is complete.
@@
Appendix C Additional Experiment Results
C.1 Skipping LEGO bricks in spatial refinement: comparing PG and PR strategies
To enhance sampling efficiency, our design allows for the skipping of a subset of LEGO bricks during sampling. During low-noise timesteps ($i.e.$, when $t$ is small), the global structure of images is already well-defined, and hence the model is desired to prioritize local details. Conversely, when images are noisier, the global-content orchestration becomes crucial to uncover global structure under high uncertainties. Therefore, for improved efficiency, we skip LEGO bricks that emphasize local details during high-noise timesteps and those that construct global strutures during low-noise timesteps.
Progressive growth (PG) and progressive refinement (PR) offer two distinct spatial refinement strategies in the LEGO model, and leveraging their unique properties to skip corresponding LEGO bricks during sampling presents a non-trivial challenge. We visualize the differences between these strategies in Figures 4 and 5. In PG, the model initially uses patch-bricks to generate patches and then employs an image-brick to re-aggregate these local features. Conversely, PR begins by utilizing the image-brick to establish a global structure and then employs local feature-oriented patch-bricks to refine details on top of it. These differences become more evident when $t$ is larger. Based on these observations, we propose that LEGO-PG can skip its top-level image-brick responsible for handling global image information when $t$ is small, while LEGO-PR skips its top-level patch-brick, which is responsible for processing $4\times 4$ local patch information, when $t$ is large.
@@ C.1 Skipping LEGO bricks in spatial refinement: comparing PG and PR strategies
Figure 6: Visualization of how FID changes as the proportion of reverse diffusion time steps, at which the top-level brick is skipped, increases from 0 to 1. The top-level brick is an image-brick whose patch size is $64\times 64$ in LEGO-S-PG and is a patch-brick whose patch size is $4\times 4$ in LEGO-S-PR. LEGO-S-PG prioritizes skipping the top-level brick at earlier time steps, while LEGO-S-PR prioritizes it at later time steps. The experiments are conducted using the LEGO-S model with eight NVIDIA A100 GPUs on the CelebA ($64\times 64$) dataset.
Figure 6: Visualization of how FID changes as the proportion of reverse diffusion time steps, at which the top-level brick is skipped, increases from 0 to 1. The top-level brick is an image-brick whose patch size is $64\times 64$ in LEGO-S-PG and is a patch-brick whose patch size is $4\times 4$ in LEGO-S-PR. LEGO-S-PG prioritizes skipping the top-level brick at earlier time steps, while LEGO-S-PR prioritizes it at later time steps. The experiments are conducted using the LEGO-S model with eight NVIDIA A100 GPUs on the CelebA ($64\times 64$) dataset.
Figure 7: Analogous results to Figure 7, conducted using the LEGO-XL model on the ImageNet ($256\times 256$) dataset with classifier-free guidance.
To validate this design, we conduct experiments to study the generation performance when LEGO bricks are skipped in different degrees. For LEGO-PG, we designate a timestep $t_{\text{break}}$ as a breakpoint. When $t>t_{\text{break}}$, no bricks are skipped, but when $t\leq t_{\text{break}}$, the top-level brick, which is the image-brick with a brick size of $64\times 64$, is skipped for the remainder of the sampling process. In contrast, for PR, we set a timestep $t_{\text{break}}$ as a breakpoint. When $t\leq T-t_{\text{break}}$, no bricks are skipped, but when $t>T-t_{\text{break}}$, the top-level brick, which is the patch-break with a patch size of $4\times 4$, is skipped for the remainder of the sampling process.
We evaluate the impact of the choice of $t_{\text{break}}$ on both FID scores and sampling time for LEGO-PG and LEGO-PR. The results, presented in Figures 7 and 7, illustrate the FID scores and the time required to generate 50,000 images. In LEGO-PG, when we skip the top-level LEGO brick, we observe a notable enhancement in sampling efficiency, primarily because this brick, responsible for processing the entire image, is computationally demanding. Nevertheless, skipping this brick during timesteps with high levels of noise also results in a trade-off with performance, as the model loses access to global content. On the other hand, for LEGO-PR, skipping the top-level patch-brick slightly impacts performance when $t_{\text{break}}$ is large, while the improvement in sampling efficiency is not as substantial as with LEGO-PG. Interestingly, when $t_{\text{break}}$ is chosen to be close to the halfway point of sampling, performance is preserved for both models, and significant time savings can be achieved during sampling.
C.2 Ablation Studies: On the effects of LEGO bricks

### v3

@@ 3.1 Main results
Training in pixel space: We begin by comparing LEGO in pixel-space training on CelebA and ImageNet with two Transformer-based state-of-the-art diffusion models (Bao et al., 2022; Peebles & Xie, 2022), which are our most relevant baselines. All images are resized to a resolution of $64\times 64$. We conduct experiments with two spatial refinement methods using LEGO, denoted as -PG for the progressive growth variant and -PR for the progressive refinement variant, as explained in Section 2.3.2. In both Tables 2 and 2, when compared to DiT, LEGO slightly increases the number of parameters due to the expansion of the input dimension in a channel-wise manner. However, in terms of FLOPs, we observe that LEGO incurs a significantly lower computational cost compared to both U-ViT and DiT. Despite maintaining such a low computational cost, LEGO still wins the competition in terms of FID performance.
@@
LEGO-XL-PR (ours, $512\times 512$) 9.01 3.99 6.39 4.87 168.92 265.75 0.76 0.86 0.63 0.49 512
@@
Training in latent space: We conduct comparisons with state-of-the-art class-conditional generative models trained on ImageNet at resolutions of $256\times 256$ and $512\times 512$, as shown in Table 3. Our results are obtained after training with 512M images (calculated as the number of iterations $\times$ batch size). When considering LEGO without classifier-free guidance (CFG), we observe that it achieves a better FID than the baseline diffusion models, striking a good balance between precision and recall. One potential explanation for this improvement is that the LEGO bricks with small patch sizes enhance patch diversity, preserving both generation fidelity and diversity.
@@
With the inclusion of CFG, we notably achieve an FID of 2.05 and the IS of 289.12 on ImageNet-$256\times 256$. It is notable that the LEGO-U variant, which incorporates the features of the -PR and -PG variants, has achieved a clearly higher IS (338.08 w. guidance and 337.85 w.o. guidance). This result significantly surpasses the baselines and demonstrates the power of multi-scale information in generative modeling. In the case of using $512\times 512$ resolution data, LEGO also achieves the best IS and produces a competitive FID with fewer iterated images. Additionally, as depicted in Figure 2, LEGO demonstrates superior convergence speed and requires significantly less training time. For qualitative validation, we provide visualizations of randomly generated samples in Figure 1.
3.2 Improving the efficiency of diffusion modeling with LEGO
The design of LEGO inherently facilitates the sampling process by generating images with selected LEGO bricks. Intuitively, at low-noise timesteps ($i.e.$, when $t$ is small), the global structure of images is already well-defined, and hence the model is desired to prioritize local details. Conversely, when images are noisier, the global-content orchestration becomes crucial to uncover global structure under high uncertainties. Therefore, for improved efficiency, we skip LEGO bricks that emphasize local details during high-noise timesteps and those that construct global structures during low-noise timesteps. To validate this design, we conduct experiments to study the generation performance when LEGO bricks are skipped in different degrees.
For LEGO-PG, we designate a timestep $t_{\text{break}}$ as a breakpoint. When $t>t_{\text{break}}$, no bricks are skipped, but when $t\leq t_{\text{break}}$, the top-level brick, which is the image-brick with a brick size of $64\times 64$, is skipped for the remainder of the sampling process. In contrast, for PR, we set a timestep $t_{\text{break}}$ as a breakpoint. When $t\leq T-t_{\text{break}}$, no bricks are skipped, but when $t>T-t_{\text{break}}$, the top-level brick, which is the patch-break with a size of $4\times 4$, is skipped for the remainder of the sampling process.
We evaluate the impact of the choice of $t_{\text{break}}$ on both FID scores and sampling time for LEGO-PG and LEGO-PR. The results in Figure 4 illustrate the FID scores and the time required to generate 50,000 images using 8 NVIDIA A100 GPUs. In LEGO-PG, when we skip the top-level LEGO brick, we observe a notable enhancement in sampling efficiency, primarily because this brick, responsible for processing the entire image, is computationally demanding. Nevertheless, skipping this brick during timesteps with high levels of noise also results in a trade-off with performance, as the model loses access to global content. On the other hand, for LEGO-PR, skipping the top-level patch-brick slightly impacts performance when $t_{\text{break}}$ is large, while the improvement in sampling efficiency is not as substantial as with LEGO-PG. Interestingly, when $t_{\text{break}}$ is chosen to be close to the halfway point of sampling, performance is preserved for both models, and significant time savings can be achieved during sampling.
Figure 4: Visualization of how FID and inference time for 50k images change as the proportion of reverse diffusion time steps, at which the top-level brick is skipped, increases from 0 to 1. LEGO-PG prioritizes skipping the top-level brick at earlier time steps, while LEGO-PR prioritizes it at later time steps. The experiments are conducted using the LEGO-XL model with eight NVIDIA A100 GPUs on the ImageNet ($256\times 256$) dataset.
3.3 Leveraging a pretrained diffusion model as a LEGO brick
Table 4: Outcomes from CIFAR10 on integrating a pre-trained model as the first brick in LEGO-PR.
Model       VP VP (frozen) + LEGO-PR   VP (unfrozen) + LEGO-PR
FID 2.5 1.92 1.88
@@ Acknowledgments
H. Zheng, Z. Wang, and M. Zhou acknowledge the support of NSF-IIS 2212418, NIH-R37 CA271186, and the NSF AI Institute for Foundations of Machine Learning (IFML).
Appendix A Discussions on related work
More efficient diffusion models
Diffusion models have garnered increasing attention from researchers seeking to enhance them in various ways. These efforts include combining variational loss with weighted denoising loss (Dhariwal & Nichol, 2021; Kingma et al., 2021; Karras et al., 2022), accelerating sampling speed using ODE solvers (Kong & Ping, 2021; San-Roman et al., 2021; Song et al., 2020; Zhang & Chen, 2022; Lu et al., 2022), and employing auxiliary models to improve training (Rombach et al., 2022; Li et al., 2022b; Zheng et al., 2022c; Pandey et al., 2022). Additionally, the iterative refinement property of diffusion models has inspired research in various domains, such as stable GAN training (Xiao et al., 2022; Wang et al., 2022), large-scale text-to-image models (Kim & Ye, 2021; Nichol et al., 2021; Gu et al., 2022; Saharia et al., 2022), and robust super-resolution and editing models (Meng et al., 2021; Choi et al., 2021; Li et al., 2022a), among others.
Backbone of diffusion models
While the downsampling and upsampling convolutional layers in U-Net allow it to effectively capture local structures, achieving a good understanding of global structures often requires a deep architecture and self-attention modules embodied in the architecture, resulting in high computational complexity during both training and generation. Instead of depending on a large number of convolutional layers, a classical approach in signal and image processing involves the extraction of local feature descriptors and their application in downstream tasks (Lowe, 2004; Mikolajczyk & Schmid, 2005; Tuytelaars & Mikolajczyk, 2008). For instance, techniques like dictionary learning and sparse coding operate on the principle that each local patch can be expressed as a sparse linear combination of dictionary atoms learned from image patches. This learned patch dictionary can then be used to enrich all overlapping patches for various image-processing tasks, such as denoising and inpainting (Aharon et al., 2006; Mairal et al., 2007; Zhou et al., 2009).
The emergence of ViT in image classification (Dosovitskiy et al., 2021), which integrates extracting feature embeddings from non-overlapping local patches with the global attention mechanism of the Transformer, marks a resurgence of this classical local-feature-based concept. ViT has not only competed effectively with state-of-the-art convolution-based networks in both image classification (Touvron et al., 2022) and generation (Esser et al., 2021), but has also been adapted to replace the U-Net in diffusion models (Yang et al., 2022; Bao et al., 2022; Peebles & Xie, 2022; Gao et al., 2023), achieving state-of-the-art image generation performance. This underscores the notion that the convolutional architecture can be supplanted by the enrichment of local features and their orchestration through global attention. ViT-based models start by enriching patch-level representations and then utilize a sequence of Transformer blocks to capture dependencies among local patches. They often demand smaller patch sizes to capture fine-grained local details and a large number of attention layers to capture global dependencies, which results in a demand of enormous data to fit. This necessitates lengthy sequence computations and deep networks to achieve optimal results, thereby incurring significant computational costs. To enhance efficiency and expedite convergence, recent works have explored hybridizing the entangled global-local structure and multi-scale properties from convolutional networks into hierarchical transformers (Liu et al., 2021; Zheng et al., 2022b; Liu et al., 2022; Yu et al., 2024).
To bridge the gap between the convolutional U-Nets and the Transformer architecture, LEGO tackles the diffusion modeling from a view of local-feature enrichment and global-content orchestration. Inspired by convolutional networks, LEGO enriches local features at both local and global scales. However, different from a series of downsampling and upsampling stages used in the U-Nets, LEGO leverages varied patch decomposition methods. This distinction results in diverse feature representations. To achieve a global-content orchestration, LEGO deploys self-attention, whose length varies depending on the ‘LEGO brick’. In some cases, attention is confined to specific image patches defined by the brick size, while in some other bricks, it extends over the entire image. This selective mechanism allows for more focused and efficient feature aggregation.
Appendix B Limitations and Future Work
@@
The current work exhibits several limitations. Firstly, we have not explored the application of LEGO bricks for text-guided image generation. Secondly, in generating class-conditional images at resolutions much higher than the training data, categories with low diversity may produce results that lack photorealism. Thirdly, our primary focus has been on stacking LEGO bricks either in a progressive-growth or progressive-refinement manner, with other stacking strategies remaining relatively unexplored. Fourthly, our selection of which LEGO bricks to skip at different time steps during generation relies on heuristic criteria, and there is room for developing more principled skipping strategies.
@@
Appendix C Experimental settings
@@
Training settings: For our LEGO model backbones, we employ EDM (Karras et al., 2022) preconditioning on pixel-space and iDDPM (Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021) on latent-space data, strictly adhering to the diffusion designs prescribed by these respective preconditioning strategies.
@@
Specifically, we employed the same suite, namely the ADM Tensorflow evaluation suite implemented in Dhariwal & Nichol (2021) [^2] https://github.com/openai/guided-diffusion/tree/main/evaluations. We also utilized their pre-extracted features for the reference batch to ensure that our evaluations are rigorous and reliable. Additionally, we tested the evaluation with scripts and reference batch provided in EDM (Karras et al., 2022)  [^3] https://github.com/nvlabs/edm. We observed that the results from both evaluation suites were consistent, showing no significant differences (up to two decimal places).
@@
During generation, for DDPM preconditioning, we use the diffusion sampler for 250 steps with a uniform stride, proposed in Nichol & Dhariwal (2021). For EDM preconditioning, we use the 2nd order Heun sampler in Karras et al. (2022). On CelebA, we sample 75 steps with the deterministic sampler. On Imagenet, we sample for 256 steps with the stochastic sampler. The stochasticity-related hyper-parameters are set as: $S_{\text{churn}}=10$, $S_{\text{min}}=0.05$, $S_{\text{max}}=20$, and $S_{\text{noise}}=1.003$.
@@
We provide an overview of the architecture configurations for the LEGO models with different capacities in Tables 5 and 6. Since Table 6 only differs from Table 5 in the brick sizes and local receptive field sizes, we will focus on explaining Table 5 in detail. Table 5 consists of three LEGO bricks, with two of them being patch-level bricks and one being an image-level brick. When using the PG-based spatial refinement approach, we stack the bricks from the bottom to the top following the order: $4\times 4\rightarrow 16\times 16\rightarrow 64\times 64$. In other words, the $4\times 4$ patch-level brick is placed at the bottom, followed by the $16\times 16$ patch-level brick, and then the $64\times 64$ image-level brick. Conversely, for the PR-based spatial refinement approach, we stack the bricks from the bottom to the top following the order: $64\times 64\rightarrow 16\times 16\rightarrow 4\times 4$. Adhering to conventional modeling techniques, we’ve outlined three distinct configurations — Small (LEGO-S), Large (LEGO-L), and XLarge (LEGO-XL) — used in our main experiments to cater to varying model capacities. Across bricks, the main differences are in token embedding and the number of attention heads, which revolves around the feature size of a local receptive field. Each brick specifies its patch size $r_{k}\times r_{k}$, local receptive field size $\ell_{k}\times\ell_{k}$, and embedding dimension (number of channels) $d_{k}$. The number of tokens (attention span) is determined as $(r_{k}/\ell_{k})^{2}$. In each brick, the configurations for LEGO-S, LEGO-L, and LEGO-XL also differ in terms of the number of attention heads and the number of DiT blocks used in the current brick to ensure scalability and adaptability across the architecture. To maintain generation coherence, we set the local-receptive-field size of the image-level brick as $2\times 2$, which effectively turns it into a DiT with a long attention span, albeit with significantly fewer layers compared to standard DiT models. Moreover, the image-level brick can be safely skipped at appropriate time steps during generation. We find the proposed configuration strikes a favorable balance between computational efficiency and performance, which is comprehensively studied in our ablation studies on the effects of LEGO bricks in Appendix D.2. During training, the LEGO patch-bricks can operate on sampled patches to further save memory and computation. In our experiments on the CelebA and ImageNet datasets, we utilize 50% and 75% of all non-overlapping patches for training patch-bricks, respectively. As noted in Section 2.2, there are instances where $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ contains missing values as some patches are not sampled in the preceding brick. In such cases, we replace these missing values with the corresponding pixels from ${\mathbf{x}}_{0}$ to ensure that the required $\hat{{\mathbf{x}}}_{0,(i,j)}^{(k-1)}$ is complete.
@@
Appendix D Additional Experiment Results
D.1 Skipping LEGO bricks in spatial refinement: comparing PG and PR strategies
Here we provide the qualitative results to support the analysis of the brick skipping mechanism of LEGO shown in Section 3.2. Progressive growth (PG) and progressive refinement (PR) offer two distinct spatial refinement strategies in the LEGO model, and leveraging their unique properties to skip corresponding LEGO bricks during sampling presents a non-trivial challenge. We visualize the differences between these strategies in Figures 5 and 6. In PG, the model initially uses patch-bricks to generate patches and then employs an image-brick to re-aggregate these local features. Conversely, PR begins by utilizing the image-brick to establish a global structure and then employs local feature-oriented patch-bricks to refine details on top of it. These differences become more evident when $t$ is larger. Based on these observations, we propose that LEGO-PG can skip its top-level image-brick responsible for handling global image information when $t$ is small, while LEGO-PR skips its top-level patch-brick, which is responsible for processing $4\times 4$ local patch information, when $t$ is large.
@@ D.1 Skipping LEGO bricks in spatial refinement: comparing PG and PR strategies
Below we show analogous results of Figure 4 on different datasets, illustrated in Figures 7-9.
Figure 7: Analogous results to Figure 4. The experiments are conducted using the LEGO-S model with eight NVIDIA A100 GPUs on the CelebA ($64\times 64$) dataset.
Figure 8: Analogous results to Figure 4, conducted using the LEGO-L model on the ImageNet ($64\times 64$) dataset with classifier-free guidance.
Figure 9: Analogous results to Figure 4, conducted using the LEGO-XL model on the ImageNet ($512\times 512$) dataset with classifier-free guidance.
D.2 Ablation Studies: On the effects of LEGO bricks